A submarine is totally submerged and moving upward. Its density is 700Kg/m^3. It was at rest 150 meters below the sea level. Find pressure on it after 3 seconds. (density of sea water is 1025 Kg/m^3).
1
Expert's answer
2012-12-13T10:21:22-0500
First we find what distance will the submarine pass in this 3 seconds upward

To do that we need to find acceleration of it.
Here it is

a=F/m = (F_a - mg)/m = (\rho_water * V * g - \rho_submarine * V * g)/\rho_submarine * V = g*(\rho_water/rho_submarine- 1) = g(1025/700 - 1) = 4.55 m/s^2

Now the distance is

s=at^2/ = 20.48 m
hence, the depth is 150-20.5 = 129.5
so, the pressure is \rho_water * g * h = 129.5*9.8*1025 = 1 300 827.5 Pa